Which of the following statements are true? Give reasons for your answers, in the form of a short proof or a counter example. (1) d^2/dx^2=(dy/dx) ^2 (2) The inverse function of y=e^3x is y=1/3lnx. (3) If f is increasing and f(x) >0 on an interval I, then g(x) = 1/f(x) is deceasing on I. (4) An equation of the tangent line to the parabola y=x^2 at (-2, 4) is y-4=2x(x+2). (5) If f is one - one onto and differentiable on R, then (f^-1) '(6) =1/f'(6)
1
Expert's answer
2020-03-16T12:07:32-0400

1)It is false. For "y=x" we have "\frac{d^2y}{dx^2}=0" and "\left(\frac{dy}{dx}\right)^2=1"

2)It is true. If "y=e^{3x}" , then "3x=\ln y" and "x=\frac{1}{3}\ln y". So inverse of "y=e^{3x}" is "y=\frac{1}{3}\ln x".

3)It is true. Take "x_1,x_2\in I", where "x_1<x_2". Since "f" is increasing on "I" and "f(x)>0" on "I", we have "0<f(x_1)<f(x_2)", so "0<\frac{1}{f(x_2)}<\frac{1}{f(x_1)}", that is "g(x_2)<g(x_1)".

4)It is false, because "y-4=2x(x+2)" is not an equation of straight line, so it is not a tangen line to curve.

5)It is not true. For "f(x)=x^3" we have "f^{-1}(x)=\sqrt[3]{x}". So "f'(x)=3x^2" and "(f^{-1})'(x)=\frac{1}{3\sqrt[3]{x^2}}".

"f'(6)=3\cdot 6^2=108" and "(f^{-1})'(6)=\frac{1}{3\sqrt[3]{6^2}}=\frac{1}{3\sqrt[3]{36}}", so "(f^{-1})'(6)\neq\frac{1}{f'(6)}"